OpenStudy (anonymous):

simplify the expression...3b^2+15b/2b^3-50b

OpenStudy (anonymous):

\[\frac{3b(b+5)}{2b(b ^{2}-25b)}= \frac{3b(b+5)}{2b(b+5)(b-5)}\] (b+5)b cancel \[\frac{3}{2(b-5)}\]